OpenStudy (anonymous):

Solve: -1 over 2 x + 10 = −x + 12

OpenStudy (anonymous):

u mean x=-1 ?

OpenStudy (anonymous):

-1 ----------------- = -x + 12 2x + 10

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

-1 ----------------- 2x + 10 = -x + 12

OpenStudy (anonymous):

no its -1 ----- x + 10 = −x + 12 2

OpenStudy (anonymous):

oh ok better now you have to combine like terms

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

I'm horrible at algebra, I don't really understand any of it ):

OpenStudy (anonymous):

lol its ok 1/2x + 10 = −x + 12

OpenStudy (anonymous):

Look this is how u work out this problem -1/2x + x = 12 - 10 I passed the Xs to one side and the numbers to the other (changing signs)

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

It kinda does, thankyou! (:

OpenStudy (anonymous):

welcomee (:

OpenStudy (anonymous):

do have multiple choice answers. I started doing the problem but I got confused at the end :)

OpenStudy (anonymous):

A. 8 B. −4 C. 1/ 2 D. 4